how can we find the slope of a line from equation of a line

Grade:12

1 Answers

Latika Leekha
9 years ago
If we have an equation of a line then we can easily find the slope from it.
If a line is in the form y = mx + b, then here ‘m’ is the slope.
Even if we have a general equation then also we can easily find the slope.
Suppose we have an equation of the line as
ax + by = c.
Then, in such a cse we can try to convert it into the form y = lx + b, so that ‘l’ will be the slope.
If a line passes through the points (x1, y1) and (x2, y2), tehn the slope is given by
m = (y2 - y1) / (x2 - x1).
These are the general cases.
You may put up a particular problem requiring the calculation of slope of a line for further clarification.
Thanks & Regards
Latika Leekha
askIITians Faculty